# Multi-Order Wavelet Derivative Transform for Deep Time Series Forecasting

### v1

@@ 3.2 WaveTS
Figure 2: The main architecture of our proposed WaveTS. For each branch where $n$ starts from 1 to $N$, first-order to $N$-th order of Wavelet Derivative Transforms (WDT) are applied to capture multi-scale representations, facilitating more effective time series modeling. LH and LL are frequency coefficients that capture representations at different scales, corresponding to micro and macro patterns, respectively.
@@ 3.2.1 Instance Normalization
Following many previous works [26, 49, 53], WaveTS uses instance normalization [19], facilitating more accurate forecasting. For the input window $\mathbf{X}_{t}=[X_{t-L+1},X_{t-L+2},\ldots,X_{t}]\in\mathbb{R}^{C\times L}$, instance normalization produces: $\operatorname{Norm}(\mathbf{X}_{t})=\left[\frac{X_{c,t}-\operatorname{Mean}_{L}(X_{c,t}^{1:L})}{\operatorname{Std}_{L}(X_{c,t}^{1:L})}\right]_{c=1}^{C},$ where $\operatorname{Mean}_{L}(X_{c,t}^{1:L})$ is the mean value along the time dimension, and $\operatorname{Std}_{L}(X_{c,t}^{1:L})$ is the corresponding standard deviation. Here, $X_{c,t}^{1:L}$ refers to the first $L$ observations of the $c$-th variate in $\mathbf{X}_{t}$.
@@ 3.2.2 WaveTS Block
Based on the proposed WDT, WaveTS consists of $N$ branches, each utilizing a WT of the total scale $K$, as shown in Figure 2. For the $n$-th branch ($n=1,2,\ldots,N$), after instance normalization, we apply an $n$-order WDT ($\mathrm{WDT}^{(n)}_{K,j}$) to the input sequence $\mathbf{X}_{t}$:
@@
where $LL_{K}^{(n)}\in\mathbb{R}^{\frac{L}{2^{K}}\times C}$ is a matrix containing the low-frequency (macro) coefficient and $\mathbf{LH}_{K}^{(n)}=\{LH_{1}^{(n)},\dots,LH_{K}^{(n)}\}$ contains multiple matrices of the $K$ high-frequency (detail) coefficients at scale $K$. Specifically, for each branch, $\mathrm{WDT}^{(n)}_{K,j}$ performs total $K$ levels of decomposition. At each decomposition level $k\in[1,K]$, the transform decomposes the current approximation coefficient $LL_{k-1}^{(n)}$ into a new approximation coefficient $LL_{k}^{(n)}$ and a detail coefficient $LH_{k}^{(n)}$. This hierarchical, tree-structured decomposition captures multi-scale information, ensuring that the final approximation coefficient $LL_{K}^{(n)}$ corresponds to the frequency band $\left[0,\frac{f_{s}}{2^{K}}\right]$, and the final detail coefficient $\mathbf{LH}_{K}^{(n)}$ covers the frequency band $\left[\frac{f_{s}}{2^{K}},f_{s}\right]$. Consequently, the full frequency spectra $\left[0,f_{s}\right]$ is comprehensively utilized. Note that the standard WT processes univariate input series rather than multivariate input, the input time series $\mathbf{X}_{t}$ is processed in a channel-independence way in each WaveTS block.
@@
After WDT-based decomposition, each branch $n$ produces one low-frequency coefficient $LL_{K}^{(n)}$ and multiple high-frequency coefficients $\mathbf{LH}_{K}^{(n)}=\{LH_{1}^{(n)},\dots,LH_{K}^{(n)}\}$. The Frequency Refinement Units, comprise multiple linear layers, subsequently refine these coefficients by expanding their length from $L$ to $L+F$ through frequency interpolation as introduced by FITS [49]. Specifically, we define a resolution factor $\alpha$ as: $\alpha=\frac{L+F}{L},$ so that each length-$L$ coefficient vector is projected to $\alpha L=L+F$ discrete points. Formally,
@@
When $K=3$, the $\mathrm{FRU}(\cdot)$ denotes a single linear layer to model $LL_{K}^{(n)}$ and the $\mathrm{FRUs}(\cdot)$ comprises 3 distinct linear layers that do not share weights to model $\mathbf{LH}_{K}^{(n)}$. Therefore, by learning these mappings for low- and high-frequency coefficients separately, the $\mathrm{FRU}(\cdot)$ effectively perform time-frequency coefficients super-resolution in a discrete format, yielding a richer spectral representation without altering the overall domain. The resulting higher-resolution frequency features can then be merged and passed to the iWDT seamlessly.
@@
We then concatenate all $N$ branches along the time dimension: $\widetilde{\mathbf{Z}}=\mathrm{Concate}\bigl{(}\mathbf{Z}_{1},\,\mathbf{Z}_{2},\,\dots,\mathbf{Z}_{N}\bigr{)}\in\mathbb{R}^{N\times(L+F)\times C}$, thus integrating the representations from multiple branches. Finally, a linear layer projects the time dimension from $N\times(L+F)$ to $L+F$, yielding the final output $\mathbf{Y}_{t}$:
@@
This design fuses multi-order wavelet derivative coefficients into one sequence while preserving the desired dimension of $C$. Finally, the output sequence $\mathbf{Y}_{t}$ is obtained using inverse instance normalization: $\operatorname{DeNorm}(\mathbf{Y}_{t})=\left[Y_{c,t}\times\operatorname{Std}_{L}(X_{c,t}^{1:L})+\operatorname{Mean}_{L}(X_{c,t}^{1:L})\right]_{1}^{C},$ where $\mathbf{Y}_{t}=[Y_{c,t-L},\ldots,Y_{c,t+F}]\in\mathbb{R}^{C\times(L+F)}$, $\operatorname{Std}_{L}(X_{c,t}^{1:L})$ and $\operatorname{Mean}_{L}(X_{c,t}^{1:L})$ originate from the input $\mathbf{X}_{t}$.
@@ 3.2.3 Training Objective
Previously, the Direct Forecast (DF) paradigm has been widely adopted in deep time series forecasting [26, 47, 65, 15]. In DF, a multi-output model $f_{\theta}$ generates predictions for the next $F$ steps in one pass: $\hat{Y}=f_{\theta}(X),$ where $X$ is the input. The L2 loss is computed by summing over all horizons: $\mathcal{L}_{DF}=\sum_{f=1}^{F}\lVert\hat{Y}_{f}-Y_{f}\rVert^{2}$, where $Y_{f}$ is the ground truth [39]. As mentioned in Section 1, the frequency features in FRL are usually processed by neural networks. However, in the simplest instantiation—e.g., a single complex-valued linear layer as in FITS [49] or a single circular convolution layer as in PaiFilter [53] and DeRiTS [7]—this mapping simplifies to a linear transformation, so the new spectra is merely a linear combination of the original. Consequently, simply reusing the previously adopted DF supervision paradigm is problematic, since it neglects the inherent presence of historical information within the future forecasting window in the frequency domain, contradicting the core principle that frequency transforms naturally preserve past information. Simply put, the output spectra already carries a “free copy" of the recent past spectra, because each frequency component is only re-scaled and phased-shifted during the linear mapping. If training focuses solely on the future window, that readily available signal is wasted, leading to suboptimal time series modeling.
Based on the aforementioned analysis, WaveTS adapts a backcasting while forecasting supervision method in training. Adding a backcasting term converts the preserved past into an extra set of labels at no additional data cost, forcing the network to reproduce what it has implicitly retained. This self-supervised consistency narrows the solution space, speeds convergence, and anchors the learned spectra-reducing the drift that appears when non-stationary series are modeled in the frequency domain [49, 3, 25]. Specifically, at each time $t$, WaveTS produces an extended output $\hat{\mathbf{Z}}_{t}\in\mathbb{R}^{C\times(L+\tau)},$ where the first $L$ steps backcast the ground-truth past series $\mathbf{X}_{t}\in\mathbb{R}^{C\times L}$ and the last $\tau$ steps forecast ground-truth $\mathbf{Y}_{t}\in\mathbb{R}^{C\times\tau}$. By concatenating $\mathbf{X}_{t}$ and $\mathbf{Y}_{t}$ along the time dimension to form $[\mathbf{X}_{t}\oplus\mathbf{Y}_{t}]\in\mathbb{R}^{C\times(L+\tau)}$, the overall L2 loss $\mathcal{L}$ is formulated as:
@@ 4.1 Experimental Setup
This paper conducts extensive experiments on 8 datasets to evaluate the performance of our proposed WaveTS in the long-term time series forecasting (LTSF) task. These datasets include Weather, Exchange, Traffic, Electricity, and ETT (four subsets). Additionally, 5 datasets are utilized for the short-term time series forecasting (STSF) task, including M4 and PEMS (four subsets). Detailed descriptions of these datasets can be found in Appendix B.1.
@@
WaveTS is compared with several SOTA models from two categories (see details in Appendix B.2), including (1) FRL-based forecasters: FITS [49], DeRiTS [7], WPMixer [31] and TexFilter [53]. (2) Other forecasters: one MLP-based model TimeMixer [41] and three representative Transformer-based models, including iTransformer [26], PatchTST [32] and FEDformer [63].
@@
Our experiments are conducted using PyTorch on a single NVIDIA RTX A6000 GPU. We utilize the ADAM [20] optimizer to minimize the backcasting while forecasting loss, as elaborated in Section 3.2.3. Furthermore, we adhere to the settings established in FITS [49] for WaveTS and the other nine baselines, considering the lookback length as a hyperparameter for grid search. This approach ensures a fair comparison, as previous literature has demonstrated that extended lookback windows optimize the performance of FRL-based forecasters [51], aligning with their training configurations. Detailed explanations of the evaluation metrics, MSE (Mean Squared Errors) and MAE (Mean Absolute Errors), are provided in Appendix B.3. The db1 basis [4] function is uniformly selected across experiments due to its real-valued property and more analysis of the wavelet basis function in the following section.
@@ 4.2 Model Comparison on Time Series Forecasting
Dataset WaveTS FITS DeRiTS WPM TexF TMix iTrans
ETTm1 0.356 0.361 0.715 0.365 0.391 0.395 0.407
ETTm2 0.244 0.252 0.321 0.264 0.285 0.276 0.288
ETTh1 0.410 0.412 0.682 0.418 0.441 0.446 0.454
ETTh2 0.332 0.337 0.435 0.354 0.383 0.365 0.383
ECL 0.160 0.172 0.293 0.175 0.172 0.183 0.178
Exchange 0.361 0.458 0.427 0.426 0.388 0.390 0.362
Traffic 0.408 0.428 0.976 0.448 0.462 0.484 0.427
Weather 0.237 0.230 0.293 0.235 0.245 0.244 0.258
Avg 0.314 0.331 0.518 0.336 0.346 0.345 0.344
Table 2: Short-term forecasting MSE on four PEMS dataset, results with MAE are provided in Table 8.
@@
Table 1 and Table 2 summarize the LTSF and STSF results, respectively, with the best in Red and the second in Blue [^1] WPM [31], TexF [53], TMix [41], iTrans[26], PatchT[32].. Only MSE are reported here due to the space limit, and Avg represents the average of the results presented. Table 1 highlights the significant superiority of WaveTS in long-term forecasting especially on datasets characterized by challenging real-world scenarios with highly intricate temporal variations, such as ECL (Electricity) and Traffic, achieving overall MSE reduction of 5.1%. In short-term forecasting, WaveTS also demonstrates an exceptional performance as summarized in Table 2. The PEMS datasets capture traffic time series data that exhibit complex fluctuations, irregular patterns, and sudden changes over time. Nonetheless, WaveTS achieves the best performance on these datasets, underscoring its effectiveness in leveraging non-stationary frequency features for time series forecasting. This paper also conduct experiments on the well-recognized M4 dataset for short-term forecasting task, the results are summarized in Table 9 in Appendix C.2.
@@ 4.3 Model Analysis
The effectiveness of the WDT is intrinsically linked to two crucial hyperparameters, i.e. the order of derivative in WDT and the scale of the wavelet transform decomposition. The former parameter dictates the number of branches, influencing the whole architecture of the WaveTS while the latter influences the hierarchical multi-scale patterns learned within each branch. Consequently, we undertake experiments to assess hyper-parameter sensitivity using the ETTh1 and Weather datasets. Figure 3 delineates the influence that these two parameters exert on achieving optimal forecasting results. Generally, the performance of the WaveTS is not sensitive to the variations in order and scale, underscoring the robustness of the WaveTS. Additionally, a larger order or scale does not guarantee better performance; therefore, we adopt a case-by-case strategy in experiments for each forecasting horizon across different datasets. We provide numerical results on more datasets in Appendix C.5. Additionally, the selection of the wavelet basis function is also crucial in wavelet-based methods; therefore, we make a detailed analysis in Appendix C.6.
@@
Ablation Study of the WDT
As previously mentioned, our proposed WaveTS, is centered on the WDT. Consequently, we conduct an ablation study to validate the necessity of the WDT for effective time series modeling. Specifically, we replace the WDT with both the Wavelet Transform (WT) and the Fourier Transform (FT), along with their corresponding inverse transformations, also maintain the instance normalizaion and subsequent Frequency Refinement Units unchanged, obtaining WaveletTS and FourierTS for comparison. Detailed description can be found in Figure 9. As demonstrated in Figure 4, the omission of the WDT leads to a significant decline in forecasting performance, thereby underscoring its essential role within the WaveTS framework. Moreover, the replacement of the WDT with the FT yields inferior performance when compared to the WT, indicating that the wavelet transform is more effective for time series analysis in FLR-based forecasters. We present additional experiments in Appendix C.7 with more results on several long-term forecasting benchmark datasets that reinforce the conclusions summarized in Figure 4.
Figure 5: Visualization of Approximation Coefficients (LL) and Detail Coefficients (LHs) in the WDT. The data are collected from the second-order derivation (order=2) at a wavelet decomposition scale of 2 (scale=2). We select the “OT” variate with the input length is 720 for each dataset.
Visualization of Coefficients in the WDT
The WDT is capable of meticulously decomposing the original time-domain series into coefficients that correspond to representations across various scales. Figure 5 visualizes the Approximation Coefficients (LL) and Detail Coefficients (LH) in the WDT of the second branch (order=2) of our proposed WaveTS. Specifically, at scale=2, we obtain two LH coefficients, denoted as LH1 and LH2, along with one LL coefficient, which is decomposed from LL1 as illustrated in Figure 2. It is evident that the LL coefficient captures the macro temporal patterns of the original series, portraying the overall trend and continuity with fewer complex fluctuations, thereby providing an effective approximation of the original series. Furthermore, LH1 and LH2 reveal high-frequency fluctuations characterized by non-periodic and stationary variations over time, thereby offering detailed and fine-grained descriptions that facilitate time series modeling. We provide more visualizations in Appendix D.2. Additionally, to complement Figure 1(i), we provide further visualizations illustrating how the proposed WDT magnifies abrupt changes in time series, enhancing better capture of explilcit temporal variations (see Figure 13 in Appendix D.3).
@@
Scalograms provide a robust framework for illustrating time-frequency characteristics within wavelet analysis, emphasizing the distribution of signal energy across various scales. In Figure 6, we present two examples at scale=5 from the ETTh1 dataset after 10 training epochs. Each row corresponds to a particular frequency band, with the lowest frequency component (LL) at the bottom and successively higher frequency details (LH1 to LH5) stacked above. Specifically, each WDT branch learns a distinct energy distribution, thus extracting different hierarchies of patterns from time series data and enhancing the model’s representational capacity. In other words, same time step may exhibit differing energy for different orders of derivation. Collectively, these multi-scale branches capture both subtle changes and global patterns, contributing to improved time series modeling. More visualizations are provided in Appendix D.4.
@@
Table 3: Efficiency summarization of WaveTS and other baselines on the ETTm1 dataset with a forecast horizon of 96. The lookback length is 360 and batch size is 64. Other configurations are set as recommended.
Model Memory Train time Inference time MSE
TimesNet [47] 21.91 GB 350.30 s 14.29 s 0.322
FEDformer [63] 7666 MB 52.40 s 10.59 s 0.325
iTransformer [26] 474 MB 5.23 s 2.69 s 0.314
PatchTST [32] 3360 MB 25.11 s 7.50 s 0.325
TexFilter [53] 450 MB 3.34 s 0.019 s 0.321
FreTS [56] 460 MB 10.88 s 0.019 s 0.343
DeRiTS [7] 396 MB 4.52 s 0.018 s 0.687
FITS [49] 394 MB 2.46 s 0.019 s 0.332
WaveTS (Ours) 380 MB 2.61 s 0.016 s 0.303
Table 3 displays the comparative results in terms of efficiency. WaveTS demonstrates superior computational efficiency than other baseline methods while maintaining accurate forecasts. Specifically, WaveTS requires only 380MB of GPU memory, takes 2.61s per epoch for training, and achieves a remarkably low inference speed of 0.016s per batch, all while obtaining an MSE of 0.303—surpassing both existing FRL-based and Transformer-based approaches in overall performance. WaveTS strikes an optimal balance between forecast accuracy and hardware usage: its efficiency gains come from the real-valued wavelet computations, which avoid the hardware-unfriendly nature of complex-valued operations common in prior FRL-based forecasters [56, 49, 63, 3, 55, 7]. This design allows WaveTS to deliver superior inference speed and low memory consumption, making it well-suited for practical deployment on edge devices and some resource-limited scenarios without sacrificing forecasting precision.

### v2

@@ 3.2 WaveTS
Figure 2: The main architecture of our proposed WaveTS. For each branch where $n$ starts from 1 to $N$, first-order to NN-th order of Wavelet Derivative Transforms (WDT) are applied to capture multi-scale representations, facilitating more effective time series modeling. LH and $L$ are frequency coefficients that capture representations at different scales, corresponding to micro and macro patterns, respectively. These coefficients are transformed back to the time domain through Inverse Wavelet Derivative Transforms (iWDT). Finally, the representations from all branches are concatenated along the time dimension and projected to obtain the final forecast.
@@ 3.2.1 Instance Normalization
Following many previous works (Liu et al., 2024c; Xu et al., 2024; Yi et al., 2024a), WaveTS uses instance normalization (Kim et al., 2021), facilitating more accurate forecasting. For the input window $\mathbf{X}_{t}=[X_{t-L+1},X_{t-L+2},\ldots,X_{t}]\in\mathbb{R}^{L\times C}$, instance normalization produces: $\operatorname{Norm}(\mathbf{X}_{t})=\left[\frac{X_{c,t}-\operatorname{Mean}_{L}(X_{c,t}^{1:L})}{\operatorname{Std}_{L}(X_{c,t}^{1:L})}\right]_{c=1}^{C},$ where $\operatorname{Mean}_{L}(X_{c,t}^{1:L})$ is the mean value along the time dimension, and $\operatorname{Std}_{L}(X_{c,t}^{1:L})$ is the corresponding standard deviation. Here, $X_{c,t}^{1:L}$ refers to the first $L$ observations of the $c$-th variate in $\mathbf{X}_{t}$.
@@ 3.2.2 WaveTS Block
[…]
@@
where $LL_{K}^{(n)}\in\mathbb{R}^{\frac{L}{2^{K}}\times C}$ is a matrix containing the low-frequency (macro) coefficient and $\mathbf{LH}_{K}^{(n)}=\{LH_{1}^{(n)},\dots,LH_{K}^{(n)}\}$ contains multiple matrices of the $K$ high-frequency (detail) coefficients at scale $K$. Specifically, for each branch, $\mathrm{WDT}^{(n)}_{K,j}$ performs total $K$ levels of decomposition. At each decomposition level $k\in[1,K]$, the transform decomposes the current approximation coefficient $LL_{k-1}^{(n)}$ into a new approximation coefficient $LL_{k}^{(n)}$ and a detail coefficient $LH_{k}^{(n)}$. This hierarchical, tree-structured decomposition captures multi-scale information, ensuring that the final approximation coefficient $LL_{K}^{(n)}$ corresponds to the frequency band $\left[0,\frac{f_{s}}{2^{K}}\right]$, and the final detail coefficient $\mathbf{LH}_{K}^{(n)}$ covers the frequency band $\left[\frac{f_{s}}{2^{K}},f_{s}\right]$. Consequently, the full frequency spectrum $\left[0,f_{s}\right]$ is comprehensively utilized. Note that the standard WT processes univariate input series rather than multivariate input, the input time series $\mathbf{X}_{t}$ is processed in a channel-independent way in each WaveTS block.
@@
After WDT-based decomposition, each branch $n$ produces one low-frequency coefficient $LL_{K}^{(n)}$ and multiple high-frequency coefficients $\mathbf{LH}_{K}^{(n)}=\{LH_{1}^{(n)},\dots,LH_{K}^{(n)}\}$. The Frequency Refinement Units comprise multiple linear layers, subsequently refine these coefficients by expanding their length from $L$ to $L+F$ through frequency interpolation as introduced by FITS (Xu et al., 2024). Specifically, we define a resolution factor $\alpha$ as: $\alpha=\frac{L+F}{L},$ so that each length-$L$ coefficient vector is projected to $\alpha L=L+F$ discrete points. Formally,
@@
When $K=3$, the $\mathrm{FRU}(\cdot)$ denotes a single linear layer to model $LL_{K}^{(n)}$ and the $\mathrm{FRUs}(\cdot)$ comprises 3 distinct linear layers that do not share weights to model $\mathbf{LH}_{K}^{(n)}$. Therefore, by learning these mappings for low- and high-frequency coefficients separately, the $\mathrm{FRU}(\cdot)$ effectively performs time-frequency coefficients super-resolution in a discrete format, yielding a richer spectral representation without altering the overall domain. The resulting higher-resolution frequency features can then be merged and passed to the iWDT seamlessly.
@@
We then concatenate all $N$ branches along the time dimension: $\widetilde{\mathbf{Z}}=\mathrm{Concat}\bigl(\mathbf{Z}_{1},\,\mathbf{Z}_{2},\,\dots,\mathbf{Z}_{N}\bigr)\in\mathbb{R}^{N\times(L+F)\times C}$, thus integrating the representations from multiple branches. Finally, a linear layer projects the time dimension from $N\times(L+F)$ to $L+F$, yielding the final output $\mathbf{Y}_{t}$:
@@
This design fuses multi-order wavelet derivative coefficients into one sequence while preserving the desired dimension of $C$. Finally, the output sequence $\mathbf{Y}_{t}$ is obtained using inverse instance normalization: $\operatorname{DeNorm}(\mathbf{Y}_{t})=\left[Y_{c,t}\times\operatorname{Std}_{L}(X_{c,t}^{1:L})+\operatorname{Mean}_{L}(X_{c,t}^{1:L})\right]_{1}^{C},$ where $\mathbf{Y}_{t}=[Y_{c,t-L},\ldots,Y_{c,t+F}]\in\mathbb{R}^{(L+F)\times C}$, $\operatorname{Std}_{L}(X_{c,t}^{1:L})$ and $\operatorname{Mean}_{L}(X_{c,t}^{1:L})$ originate from the input $\mathbf{X}_{t}$.
@@ 3.2.3 Training Objective
Previously, the Direct Forecast (DF) paradigm has been widely adopted in deep time series forecasting (Liu et al., 2024c; Wu et al., 2023; Zhou et al., 2024; Huang et al., 2023). In DF, a multi‐output model $f_{\theta}$ generates predictions for the next $F$ steps in one pass: $\hat{Y}=f_{\theta}(X),$ where $X$ is the input. The L2 loss is computed by summing over all horizons: $\mathcal{L}_{DF}=\sum_{f=1}^{F}\lVert\hat{Y}_{f}-Y_{f}\rVert^{2}$, where $Y_{f}$ is the ground truth (Wang et al., 2025a). As mentioned in Section 1, the frequency features in FRL are usually processed by neural networks (we illustrate the FRL paradigm for better understanding in Appendix A). However, in the simplest instantiation, e.g., a single complex-valued linear layer as in FITS (Xu et al., 2024) or a single circular convolution layer as in PaiFilter (Yi et al., 2024a) and DeRiTS (Fan et al., 2014), this mapping simplifies to a linear transformation, so the new spectrum is merely a linear combination of the original. Consequently, simply reusing the previously adopted DF supervision paradigm is problematic, since it neglects the inherent presence of historical information within the future forecasting window in the frequency domain, contradicting the core principle that frequency transforms naturally preserve past information. Simply put, the output spectrum already carries a “free copy” of the recent past spectrum, because each frequency component is only re-scaled and phased-shifted during the linear mapping. If training focuses solely on the future window, that readily available signal is wasted, leading to suboptimal time series modeling.
Based on the aforementioned analysis, WaveTS adopts a backcasting while forecasting supervision method in training. Adding a backcasting term converts the preserved past into an extra set of labels at no additional data cost, forcing the network to reproduce what it has implicitly retained. This self-supervised consistency narrows the solution space, speeds convergence, and anchors the learned spectrum, reducing the drift that appears when non-stationary series are modeled in the frequency domain (Xu et al., 2024; Cao et al., 2020; Liu et al., 2024b). Specifically, at each time $t$, WaveTS produces an extended output $\hat{\mathbf{Z}}_{t}\in\mathbb{R}^{(L+\tau)\times C},$ where the first $L$ steps backcast the ground-truth past series $\mathbf{X}_{t}\in\mathbb{R}^{L\times C}$ and the last $\tau$ steps forecast ground-truth $\mathbf{Y}_{t}\in\mathbb{R}^{\tau\times C}$. By concatenating $\mathbf{X}_{t}$ and $\mathbf{Y}_{t}$ along the time dimension to form $\mathbf{X}_{t}\oplus\mathbf{Y}_{t}]\in\mathbb{R}^{(L+\tau)\times C}$, the overall L2 loss $\mathcal{L}$ is formulated as:
@@ 4.1 Experimental Setup
This paper conducts extensive experiments on 8 datasets to evaluate the performance of our proposed WaveTS in the long-term time series forecasting (LTSF) task. These datasets include Weather, Exchange, Traffic, Electricity (ECL), and ETT (four subsets). Additionally, 5 datasets are utilized for the short-term time series forecasting (STSF) task, including M4 and PEMS (four subsets). Detailed descriptions of these datasets can be found in Appendix C.1.
@@
WaveTS is compared with several SOTA models from two categories (see details in Appendix C.2), including (1) FRL-based forecasters: FITS (Xu et al., 2024), DeRiTS (Fan et al., 2014), WPMixer (Murad et al., 2025) and TexFilter (Yi et al., 2024a). (2) Other forecasters: one hybrid model TimeMixer++ (Wang et al., 2025b), two representative Transformer-based models, including iTransformer (Liu et al., 2024c), PatchTST (Nie et al., 2023) and one MLP-based model DLinear (Zeng et al., 2023). We follow the official configurations when reproducing the results.
@@
Our experiments are conducted using PyTorch on a single NVIDIA RTX A6000 GPU. We utilize the ADAM (Kingma and Ba, 2014) optimizer to minimize the backcasting while forecasting loss, as elaborated in Section 3.2.3. Furthermore, we adhere to the settings established in FITS (Xu et al., 2024) for WaveTS and the other nine baselines, considering the lookback length as a hyperparameter for grid search. This approach ensures a fair comparison, as previous literature has demonstrated that extended lookback windows optimize the performance of FRL-based forecasters (Yang et al., 2024), aligning with their training configurations. Detailed explanations of the evaluation metrics, MSE (Mean Squared Errors) and MAE (Mean Absolute Errors), are provided in Appendix C.3. The db1 basis (Daubechies, 1988) function is uniformly selected across experiments due to its real-valued property and more analysis of the wavelet basis function in Appendix D.8.
@@ 4.2 Model Comparison on Time Series Forecasting
Dataset WaveTS FITS DeRiTS WPM TexF TMix+ iTrans
ETTm1 0.356 0.361 0.715 0.365 0.391 0.377 0.407
ETTm2 0.244 0.252 0.321 0.264 0.285 0.269 0.288
ETTh1 0.410 0.412 0.682 0.418 0.441 0.419 0.454
ETTh2 0.332 0.337 0.435 0.354 0.383 0.339 0.383
ECL 0.160 0.172 0.293 0.175 0.172 0.165 0.178
Exchange 0.361 0.458 0.427 0.426 0.388 0.357 0.362
Traffic 0.408 0.428 0.976 0.448 0.462 0.416 0.427
Weather 0.237 0.230 0.293 0.235 0.245 0.226 0.258
Avg 0.314 0.331 0.518 0.336 0.346 0.320 0.344
Table 2: Short-term forecasting MSE on four PEMS datasets, results with MAE are provided in Table 8.
@@
Table 1 and Table 2 summarize the LTSF and STSF results, respectively, with the best in Red and the second in Blue [^1] WPM (Murad et al., 2025), TexF (Yi et al., 2024a), TMix+ (Wang et al., 2025b), iTrans (Liu et al., 2024c), PatchT (Nie et al., 2023).. Only MSE are reported here due to the space limit, and Avg represents the average of the results presented. Table 1 highlights the significant superiority of WaveTS in long-term forecasting especially on datasets characterized by challenging real-world scenarios with highly intricate temporal variations, such as ECL (Electricity) and Traffic, achieving overall MSE reduction of 5.1%. In short-term forecasting, WaveTS also demonstrates exceptional performance as summarized in Table 2. The PEMS datasets capture traffic time series data that exhibit complex fluctuations, irregular patterns, and sudden changes over time. Nonetheless, WaveTS achieves the best performance on these datasets, underscoring its effectiveness in leveraging non-stationary frequency features for time series forecasting. This paper also conducts experiments on the well-recognized M4 dataset for short-term forecasting task, the results are summarized in Table 9 in Appendix D.2.
@@ 4.3 Model Analysis
The effectiveness of the WDT is intrinsically linked to two crucial hyperparameters, i.e. the order of derivative in WDT and the scale of the wavelet transform decomposition. The former parameter dictates the number of branches, influencing the whole architecture of the WaveTS while the latter influences the hierarchical multi-scale patterns learned within each branch. Consequently, we undertake experiments to assess hyper-parameter sensitivity using the ETTh1 and Weather datasets. Figure 3 delineates the influence that these two parameters exert on achieving optimal forecasting results. Generally, the performance of the WaveTS is not sensitive to the variations in order and scale, underscoring the robustness of the WaveTS. Additionally, a larger order or scale does not guarantee better performance; therefore, we adopt a case-by-case strategy in experiments for each forecasting horizon across different datasets. We provide numerical results on more datasets in Appendix D.5. In Appendix D.6, we also justify the principle for selecting hyperparameters with more experiments. In Appendix D.7, we quantitatively analyze the contribution of each WaveTS block to the final prediction and identify the order and scale configurations most suitable for different dataset types. Additionally, the selection of the wavelet basis function is also crucial in wavelet-based methods; therefore, we make a detailed analysis in Appendix D.8.
Ablation Study of the WDT
@@
As previously mentioned, our proposed WaveTS, is centered on the WDT. Consequently, we conduct an ablation study to validate the necessity of the WDT for effective time series modeling. Specifically, we replace the WDT with both the Wavelet Transform (WT) and the Fourier Transform (FT), along with their corresponding inverse transformations, also maintain the instance normalization and subsequent Frequency Refinement Units unchanged, obtaining WaveletTS and FourierTS for comparison. Detailed description can be found in Figure 9 in Appendix D.9. As demonstrated in Figure 4, the omission of the WDT leads to a significant decline in forecasting performance, thereby underscoring its essential role within the WaveTS framework. Moreover, the replacement of the WDT with the FT yields inferior performance when compared to the WT, indicating that the wavelet transform is more effective for time series analysis in FRL-based forecasters. We present additional experiments in Appendix D.9 with more results on several long-term forecasting benchmark datasets that reinforce the conclusions summarized in Figure 4.
Visualizations of the Abrupt Changes Amplification
As illustrated in Figure 5, across all four datasets, the highlighted windows show that the magnitude of the wavelet derived coefficients surges precisely when the original series experiences a swift trend break and volatility jump. This behavior is consistent with the analytical role of the derivative operation as a local high-pass filter: it suppresses slow-moving components while accentuating rapid derivatives. By contrast, the standard WT coefficients remain comparatively smooth, indicating limited sensitivity to such localized shocks. The repetition of this pattern on time series collected from various domain demonstrates that the WDT generalizes beyond a single domain, providing a robust mechanism for exposing change points that would otherwise be under-represented in frequency representations.
Figure 5: Additional visual evidence that the Wavelet Derived Coefficients (red) produced by the proposed WDT amplify abrupt regime shifts more clearly than the standard Wavelet Coefficients (blue).
@@
Scalograms provide a robust framework for illustrating time-frequency characteristics within wavelet analysis, emphasizing the distribution of signal energy across various scales. In Figure 6, we present two examples at scale=5 from the ETTh1 dataset after 10 training epochs. Each row corresponds to a particular frequency band, with the lowest frequency component (LL) at the bottom and successively higher frequency details (LH1 to LH5) stacked above. Specifically, each WDT branch learns a distinct energy distribution, thus extracting different hierarchies of patterns from time series data and enhancing the model’s representational capacity. In other words, same time step may exhibit differing energy levels across different orders of derivation. Collectively, these multi-scale branches capture both subtle changes and global patterns, contributing to improved time series modeling. More visualizations are provided in Appendix E.3.
@@
Table 3: Efficiency summary of WaveTS and other baselines on the ETTm1 dataset with a forecast horizon of 96. The lookback length is 360 and batch size is 64. Other configurations are set as recommended.
Model Memory Train Infer. MSE
FEDformer (Zhou et al., 2022b) 7666 MB 52.40s 0.152s 0.325
iTransformer (Liu et al., 2024c) 474 MB 5.23s 0.053s 0.314
PatchTST (Nie et al., 2023) 3360 MB 25.11s 0.076s 0.325
TimeMixer++ (Wang et al., 2025b) 17506 MB 30.21s 0.104s 0.310
TexFilter (Yi et al., 2024a) 450 MB 3.34s 0.021s 0.321
FreTS (Yi et al., 2023b) 460 MB 10.88s 0.024s 0.343
DeRiTS (Fan et al., 2014) 396 MB 4.52s 0.018s 0.687
FITS (Xu et al., 2024) 394 MB 2.46s 0.019s 0.332
WaveTS (Ours) 380 MB 2.61s 0.015s 0.303
Table 3 displays the comparative results in terms of efficiency. “Train” and “Infer.” refer to training time per epoch and inference time per batch, respectively. WaveTS demonstrates superior computational efficiency than other baseline methods while maintaining accurate forecasts. Specifically, WaveTS requires only 380MB of GPU memory, takes 2.61s per epoch for training, and achieves a remarkably low inference time of 0.015s per batch, all while obtaining an MSE of 0.303, surpassing both existing FRL-based and Transformer-based approaches in overall performance. WaveTS strikes an optimal balance between forecast accuracy and hardware usage: its efficiency gains come from the real-valued wavelet computations, which avoid the hardware-unfriendly nature of complex-valued operations common in prior FRL-based forecasters (Yi et al., 2023b; Xu et al., 2024; Zhou et al., 2022b; Cao et al., 2020; Yi et al., 2023a; Fan et al., 2014). This design allows WaveTS to deliver superior inference speed and low memory consumption, making it well-suited for practical deployment on edge devices and some resource-limited scenarios without sacrificing forecasting precision.